The class average on a math test was 82 and there are 8 students in the class. If seven of the scores are 80, 90, 78, 100, 76, 74, and 82, what must the 8th score be to produce the average of 82?
Answer by richwmiller(17219)

(80+90+78+100+76+74+82+x)/8=82
(580+x)/8=82
580+x=656
x=76